Fermi National Laboratory
 

Operation Summary for September 17, 2010 - September 20, 2010

 

The Integrated Luminosity for the period from 9/13/10 to 9/20/10 was 60.37 pb-1
NuMI reported receiving 5.2E18 protons on target during this same period.

Friday Sept 17

The day shift began with Operations monitoring the Tevatron (TeV) store 8093 (95.22E30) and the Recycler stash (276.4E10), with the Antiproton Source (Pbar) stacking (21.9E10), with the Linac, Booster, the and Main Injector (MI) operating normally, with the NuMI experiment off for target work, with MiniBooNE and the Meson MTest T992 and T995 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations terminated store 8093 at 9:40 AM. Cryo system technicians began work on the TeV sector C2 wet engine. NuMI personnel performed target scans and then accessed
their hall to pull the target out for inspection and to repair a helium leak, they expect to be off for eight days. Operations put the TeV into HEP shot setup at 1:20 PM and established store 8096 at 2:30 PM with an average initial luminosity (AIL) of 263E30. Linac personnel began their power amplifier (PA) change on a RF station (LRF5).

During the evening shift, Linac personnel reported at 5:20 PM that they had completed their PA change. Operations established beam to Pbar, MiniBooNE, and MTest. Operations disabled the MTest beam at 6 PM. At 7:15 PM, a Pbar extraction kicker (D:EKIK) tripped off due to a failed cable that was arcing. Pbar personnel came in and replaced the cable. Stacking resumed at 10:05 PM.

Saturday Sept 18

The midnight shift began with Operations monitoring store 8096 (82.84E30) and the stash (193.8E10), with Pbar stacking (14.0E10), and with MiniBooNE taking beam.

During the midnight shift, at 1:05 AM, the Linac vacuum valves at the 400 MeV area Lambertson
and the dump closed. Operations called in Mechanical Support personnel and Linac personnel to investigate. They determined that a shorted cable for a vacuum gauge (BD-2) was the cause. They temporarily bypassed the gauge interlock and opened the vacuum valves. Linac resumed sending beam at 4:10 AM. Operations reported that a MI injection kicker (I:KPS1C) had tripped off numerous times on a prefire indication. Operations contacted the system engineer and he had operators lower the kicker power supply voltage and increase the voltage on kickers KPS1A and KPS1B. LRF1 tripped off several times due to a permanent inhibit. Operations lowered the pulse-width on LRF1 for a study and then returned the station to normal operation.

During the day shift, Operations established beam to MTest at 8 AM. Operations terminated store 8096 at 8:50 AM, put the TeV into shot setup, and established store 8097 at 10:20 AM with an AIL of 275E30.

During the evening shift, Operations disabled the MTest beam at 6 PM. Operations terminated store 8097 at 11:55 PM.

Sunday Sept 19

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE taking beam.

During the midnight shift, Operations put the TeV into shot setup at 12:05 AM and established store 8098 at 1:10 AM with an AIL of 309E30.

During the day shift, Operations terminated store 8098 at 3 PM, put the TeV into shot setup and established store 8099 at 10:05 AM with an AIL of 289E30. Operations reported that the Boostergradient magnet power supply (GMPS) tripped off at 7:45 PM.

During the evening shift, Booster personnel came in and investigated the GMPS control system; they found a faulty control cable and repaired it. Booster resumed sending beam at 10:30 PM. At 11:50 PM, the H- Source tripped off.

Monday Sept 20

The midnight shift began with Operations monitoring store 8099 and the stash, and with Pbar and MiniBooNE off due to the H- Source problem.

During the midnight shift, operators had the H- Source back online by 12:05 AM. Operations established beam to MTest at 6:50 AM.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run coordinator will allow MI personnel to conduct a coalescing study.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/20/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory